OpenStudy (anonymous):

How long is a leg a of a right triangle whose other leg b measures 2x units, and whose hypotenuse c measures 4x units (where x 0) ?

OpenStudy (anonymous):

this is sq root ((2x)^2 + (4x)^2))

OpenStudy (anonymous):

oopps - sorry it is sq root((4x)^2 - (2x)^2))

OpenStudy (anonymous):

\[(4x)^{2}-(2x)^{2}\]

OpenStudy (anonymous):

this comes to sq root12 x^2

OpenStudy (anonymous):

\[16x^{2}-4x ^{2}=12x ^{2}\]\[\sqrt{12x ^{2}}=2x \sqrt{3}\]

OpenStudy (anonymous):

\[\frac{\sqrt{3}}{2}x \]

OpenStudy (anonymous):

Thank you so much! I havent had math in a year and I'm trying to review some math

OpenStudy (anonymous):

If I deserve medal then give me one.

OpenStudy (anonymous):

How do I do that..

OpenStudy (anonymous):

next to my name click give medal

OpenStudy (anonymous):

what u see ,give medal or good answer?

OpenStudy (anonymous):

\[4x \text{Cos}[\text{ArcSin}[(2x)/(4x)]] =2 \sqrt{3} x \] I forgot to multiply by 4x

OpenStudy (anonymous):

thank

myininaya (myininaya):

\[2x \sqrt{4-x^2}\]

myininaya (myininaya):

its a right triangle so we have a^2+b^2=c^2 c=4x b=2x a=? a^2=c^2-b^2 a=sqrt{c^2-b^2} a=sqrt{(4x)^2-(2x)^2} a=sqrt{16x^2-4x^2} a=sqrt{4x^2}sqrt{4-1} a=2xsqrt{3} oh i made a mistake lol i guess it helps when u work it out